OpenStudy (anonymous):

If the density of gasoline is 0.66g/mL,how many kilocalories of energy are obtained from 2.8gal of gasoline? When 1.0g of gasoline burns, 11 500kcal of energy are given off.

OpenStudy (xishem):

\[2.8gal(gasoline)*\frac{3.79L}{1gal}*\frac{1000mL}{1L}=10612mL(gas)\]\[D=\frac{m}{V} \rightarrow m=DV\]\[m=(0.66g*mL^{-1})(10612mL)=7003.92g(gas)\]\[7003.92g(gas)*\frac{11,500kcal}{1.0g(gas)}=80,545,080kcal\]We need to report it to 2SFs...\[8.1*10^7kcal\]

OpenStudy (anonymous):

omg thanks sooo much